OpenStudy (anonymous):

f(x)=4x2-5x (the 2 is for the square) to get answers for f(0)= f(-1)= f(2)=

OpenStudy (anonymous):

they are asking you to plug values in for x, f(0)=4(0)^2 - 5(0) and so on for each value

OpenStudy (anonymous):

i know that but i cant seem to find the answers except for the first one

OpenStudy (anonymous):

the setup is the same f(-1)=4(-1)^2 - 5(-1), just watch your signs

myininaya (myininaya):

=4(1)+5=4+5=9

OpenStudy (anonymous):

f(2) = 4*2^2 - 5*2 = 6